OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. 3x – 5y + 2z = 7 4x + 9y – z = 2 x – 3y + 3z = 7

hero (hero):

Nobody like triple variables

OpenStudy (anonymous):

I love triple variables

OpenStudy (anonymous):

first, I would move the z terms so they are the first ones (because that's how I roll)

OpenStudy (anonymous):

then I would use Gaussian elimination

OpenStudy (anonymous):

3x – 5y + 2z = 7 [1] 8x + 18y – 2z = 4 [2] *2 =========== 11x+13y =11 [1]' 12x + 27y – 3z = 2 [1]*3 x – 3y + 3z = 7 [3] ============ 13x+24y =9 [3]' The 2x2 system would be: 11x+13y =11 [1]' 13x+24y =9 [3]' Let's hope I didn't make an arith. error!

OpenStudy (anonymous):

where I have [1]*3 that is supposed to read [2]*3

OpenStudy (anonymous):

z + 1/3x -y = 7/3 x + (24/13)y = 1 y = 0

OpenStudy (anonymous):

y = 0, x = 1, z = 2

OpenStudy (anonymous):

I did have an error--in my second linear combination. these problems are not best solved on the computer! 12x + 27y – 3z = 6 [2]*3 x – 3y + 3z = 7 [3] ============ 13x+24y =13 [3]' The 2x2 system would be: 11x+13y =11 [1]' 13x+24y =13 [3]'

OpenStudy (anonymous):

this seems to be ok b/c it checks with agd...o's answers